At a company fish fry, ½ in attendance are employees. Employees’ spouses are 1/3 of the attendance. What is the percentage of the people in attendance who are not employees or employee spouses?
Answer by AnlytcPhil(1806)

At a company fish fry, 1/2 in attendance are employees.
Employees’ spouses are 1/3 of the attendance. What is 
the percentage of the people in attendance who are not 
employees or employee's spouses? 

One half of X are employees.
One third of X are employee's spouses.

We add one half PLUS one third:

1/2 + 1/3 = 3/6 + 2/6 = 5/6 

Then we subtract that fraction from one whole, or 1,
to see what fraction is left.  That is, we say

1 minus 5/6

which is

1 - 5/6

Now write 1 as 6/6

6/6 - 5/6

We get 1/6

So 1/6 is left.  Now we need to make that into
a percent by multiplying it by 100 and tacking on a "%"

1/6 × 100

1/6 × 100/1

100/6

50/3

16 2/3 %

So 16 2/3 % of the people in 
attendance are neither employees
nor employee's spouses.


Now let's check:

1/2 are employees.  That's 50%
1/3 are wmployee's spouses. That's 33 1/3%
15 2/3 % are neither employees nor employee's spouses.

Add them up

50     %
33 1/3 %
16 2/3 %
--------
99 3/3 %

And 99 3/3% = 100%

Edwin